OpenStudy (anonymous):

how do i find the sin 11pi/2 without using a calculator?

OpenStudy (anonymous):

thats a fix value. sin (pi/2) = 1

OpenStudy (anonymous):

\[\frac{11 \pi}{2}\] on the unit circle and look at the second coordinate

OpenStudy (anonymous):

i get it that it = 990 degrees but how do i get it in point form

OpenStudy (anonymous):

\[\sin(\frac{11\pi}{2})=\sin(\frac{10\pi}{2}+\frac{\pi}{2})\]

OpenStudy (anonymous):

if you like you can a) count b) subtract of 2pi repeatedly until you get a number between 0 and 2 pi

OpenStudy (amistre64):

draw a circle; and place an angle that sweeps out 11pi/2 ; and measure how far its terminal end point is from the x axis :)

OpenStudy (anonymous):

Oh.. It's 11pi/2. Sorry i misread the question. (11/2)pi = 5pi + pi/2 5pi = two and a half turn around the trignometric circle. if you sum more pi/2 then you're in (3/2)pi = -1

OpenStudy (anonymous):

\[\sin(5\pi+\frac{\pi}{2})\] this is as far as i can go my teacher always said to chance it

OpenStudy (anonymous):

in any case the answer is NOT 1

OpenStudy (anonymous):

lets do it the simple donkey way. make life easy

OpenStudy (anonymous):

\[\sin x = \sum_{n=1}^{\infty} (-1/(2n+1)!)*x^{2n+1}\]

OpenStudy (anonymous):

\[\frac{11}{2}=5\tfrac{1}{2}\] subtract 4 get \[1\tfrac{1}{2}=\frac{3}{2}\]

OpenStudy (anonymous):

sin(5pi)cos(pi/2)+sin(pi/2)cos(5pi) 0*0+1*(-1) -1

OpenStudy (anonymous):

so just locate \[]\frac{3\pi}{2}\] on the unit circle. you will see that the second coordinate is -1 (not 1) and that is your answer

OpenStudy (anonymous):

i just plugged it into the formula i seen online

OpenStudy (anonymous):

this is all way way way too much work. give the guy (gal) a break

OpenStudy (anonymous):

thanks everyone i get it now

OpenStudy (anonymous):

really this is very easy and straight forward

OpenStudy (anonymous):

wait satelite why did you subtract 4?

OpenStudy (anonymous):

no formula, not infinite sum lordamercy. just look at \[\frac{3\pi}{2}\] and see where you are. finito

OpenStudy (anonymous):

11pi/2=10pi/2+1pi/2 right? 5pi+pi/2

hero (hero):

Draw a Unit Circle, label, pi, pi/2, 3pi/2, and 2pi....Then start with pi/2 and keep going around the unit circle until you get to 11pi/2. one pi/2....2pi/2...3pi/2....etc.

OpenStudy (anonymous):

@chelsea i subtract 4 because i am working with \[\frac{11}{2}\] instead of \[\frac{11\pi}{2}\] sometimes the pi throws you off

OpenStudy (anonymous):

so what i was really doing was subtracting off \[2\pi\] as many times as i could to end up between 0 and 2 pi

hero (hero):

My method requires no adding, subtracting, multiply, dividing or any of that complicated stuff, just the ability to count and go around a unit circle.

OpenStudy (anonymous):

for example if i see \[\frac{25\pi}{2}\] i think to myself \[\frac{25}{2}=12\tfrac{1}{2}\] and 2 goes in to 12 evenly so this is the same as \[\frac{1}{2}\] i.e. \[\frac{25\pi}{2}\] is coterminal with \[\frac{\pi}{2}\] and i look there.

OpenStudy (anonymous):

@hero of course you are right and i frequently count as well. but suppose you had \[\frac{25\pi}{2}\] are you really going to count?

OpenStudy (anonymous):

counting to eleven is ok but i have no desire to count to twenty five going around and around and around

OpenStudy (anonymous):

i mean i have done exactly what you wrote many times. one pi over two two pi over two three pi over two four pi over two etc, but at some point you do not want to do this

OpenStudy (anonymous):

dude... if it's even than its an entire turn around the circle. which means you're in position 0 again. or 2pi if you want. if it's odd than you're in half turn of only pi

OpenStudy (anonymous):

oooooh i get it so just simplify and it will be coterminal ?

OpenStudy (anonymous):

right exactly. you want to end up between 0 and 2 pi

hero (hero):

You can use other methods to get there quicker.....12pi/2 = 6 pi.....24/2 = 12pi which are all multiples of 2pi, you just count one more, and you have the eqivalent of 25pi/2 which is pi/2

OpenStudy (anonymous):

by subtracting off two pi as many times as you can. which is the same as dividing by 2 and taking the remainder

OpenStudy (anonymous):

which is what hero is saying if i understand him correctly

OpenStudy (anonymous):

3pi/2 around the circle then pi/2+3pi/2 to end up at same spot pi/2+(pi/2+3pi/2) to end again at same point ;;;;; 8*pi/2+3pi/2

hero (hero):

No....I'm counting....

hero (hero):

Either way, the unit circle method is simpler

OpenStudy (anonymous):

actually you divided 12 by 2 just as i did. it is the same

hero (hero):

Ah, there's a subtle difference

OpenStudy (anonymous):

you can call it counting but you wrote \[\frac{12\pi}{2}=6\pi\] and i said 2 goes in to twelve evenly, this is the same

OpenStudy (anonymous):

believe me i am not trying to say that you are wrong, just that we are really doing the same thing

hero (hero):

The difference is, I don't keep subtracting 2pi, I only started with 12pi/2 for explanation purposes....I could have simply started with 24pi/2 = 12pi which is a multiple of 2pi

hero (hero):

and then simple count 1 more and I have pi/2

OpenStudy (anonymous):

hey satelite what did u do with that extra pi/2 left over again?

OpenStudy (anonymous):

yes and i maintain that they are the same procedure. saying the remainder when i divide 25 by 2 is 1 is the same as saying i subtract 2 12 times leaving 1. of course i am not actually subtracting 2 12 times, i am dividing

hero (hero):

My method works for those who are not adept and subtracting and dividing...that's the difference

OpenStudy (anonymous):

@chelsea you locate \[\frac{\pi}{2}\] on the unit circle, and look at the second coordinate of you want sine and first if you want cosine. and excellent cheat sheet is here http://tutorial.math.lamar.edu/cheat_table.aspx

OpenStudy (anonymous):

look at the trig cheat sheet. it has a nice unit circle on it

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

unit circle on last page it has it all mapped out for you

hero (hero):

You don't add, you don't subtract, you don't multiply, you don't divide. You just simply "count". It's the same with those sensor activated toilet flushers. You don't touch the lever, you just simply dump and go

OpenStudy (anonymous):

@hero what is coterminal with \[\frac{127\pi}{4}\] and how would you find it without dividing?

OpenStudy (anonymous):

ooh one last thing so what i locate is the remainder for all other problems>?

OpenStudy (anonymous):

i do like the flush analogy

hero (hero):

You got me on that one....I was referring to stuff under 30pi/2 or 30pi/4....anything like that and you'll at least have to divide

hero (hero):

find the nearest muliple of 4

hero (hero):

Which is 31pi, then you're left with 3pi/4